OpenStudy (anonymous):

Consider the functions f(x) = 3x2 + 1 and g(x) = x + 3. What is the value of f[g(−1)]? 7 37 13 12

OpenStudy (anonymous):

what is f(x) eq?

OpenStudy (anonymous):

f(g(-1)= f(-1+3)=f(2)=3(2)^2+1=3(4)+1=12+1=13

OpenStudy (anonymous):

C???

OpenStudy (anonymous):

it says to take g function and evaluate at -1 so -1+3=2 so then take f function and evaluate at 2 so \[3(2)^{2}+1\]

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

ya i guess

OpenStudy (anonymous):

yes or no not ya?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

good job